# Towards Realistic Scene Generation with LiDAR Diffusion Models

### v1

@@ Abstract
Diffusion models (DMs) excel in photo-realistic image synthesis, but their adaptation to LiDAR scene generation poses a substantial hurdle. This is primarily because DMs operating in the point space struggle to preserve the curve-like patterns and 3D geometry of LiDAR scenes, which consumes much of their representation power. In this paper, we propose LiDAR Diffusion Models (LiDMs) to generate LiDAR-realistic scenes from a latent space tailored to capture the realism of LiDAR scenes by incorporating geometric priors into the learning pipeline. Our method targets three major desiderata: pattern realism, geometry realism, and object realism. Specifically, we introduce curve-wise compression to simulate real-world LiDAR patterns, point-wise coordinate supervision to learn scene geometry, and patch-wise encoding for a full 3D object context. With these three core designs, our method achieves competitive performance on unconditional LiDAR generation in 64-beam scenario and state of the art on conditional LiDAR generation, while maintaining high efficiency compared to point-based DMs (up to 107$\times$ faster). Furthermore, by compressing LiDAR scenes into a latent space, we enable the controllability of DMs with various conditions such as semantic maps, camera views, and text prompts. Our code and pretrained weights are available at https://github.com/hancyran/LiDAR-Diffusion.
Figure 1: Our method (LiDM) steps towards LiDAR-realistic scene generation by preserving curve-like structures and objects with greater resemblance to real-world data (Reference), and marks a milestone for conditional LiDAR scene generation from different input modalities.
@@ 1 Introduction
  • Our method achieves state-of-the-art on conditional scene synthesis under 64-beam scenario, while realizing a $\times$107 speedup compared to point-based DMs.
@@ 2 Related Work
Given the larger-scale and complex scenes, treating LiDAR point clouds as a point cloud generation task akin to hand-crafted 3D models encounters difficulties. In [7], the authors explore the possibilities of generative models in LiDAR scenes by providing two solutions, LiDARVAE and LiDARGAN. [73] further introduces a point-based DM, LiDARGen, to generate satisfactory samples of range images as LiDAR scenes. However, these aforementioned methods still fail to generate LiDAR-realistic scenes as they may overlook the curve-like structures and geometric details inherent in LiDAR data.
@@ 4.2 Unconditional LiDAR Diffusion
Perceptual Statistical Method FRID $\downarrow$ FSVD $\downarrow$ FPVD $\downarrow$ JSD $\downarrow$ MMD $\downarrow$ ($10^{-4}$) Noise 3277 497.1 336.2 0.360 32.09 LiDARGAN [7] 1222 183.4 168.1 0.272 4.74 LiDARVAE [7] 199.1 129.9 105.8 0.237 7.07 LiDARGen [73] (1160s) 129.1 39.2 33.4 0.188 2.88 LiDARGen [73] (50s) 2051 480.6 400.7 0.506 9.91 LDM [51] (50s) 199.5 70.7 61.9 0.236 5.06 LiDM (ours, 50s) 158.8 53.7 42.7 0.213 4.46 $\Delta$ Improv. 20.4% 24.0% 31.0% 9.7% 11.9%
Table 1: Comparison of unconditional LiDAR scene generation with recent state-of-the-art methods. We conduct experiments on 64-beam (i.e., KITTI-360 [37]) data.$\downarrow$” indicates that lower values are better. $N$-s refers to $N$ sampling steps during inference. $\Delta$ Improv. is the relative improvement of our method compared to the baseline of Latent Diffusion (LDM) [51], with the same number of diffusion steps. Note that, the diffusion process of LiDARGen [73] has 232 levels and 5 iterations in each level (i.e., 1160 steps in total).   denotes baseline results, while   denotes our results.
To verify the effectiveness of our method, we train unconditional LiDMs on 64-beam data from KITTI-360 [37]. We evaluate the performance of each method using FRID (range image), FSVD (sparse volume), and FPVD (point-based volume), as described in Sec. 3.6. Together, these three metrics enable the perceptual evaluation to assess the quality of LiDAR scenes generated with different methods.
As shown in Table 1, with a very limited number of sampling steps (i.e., 50), we establish a new state of the art for almost all considered metrics. Specifically, within 50 sampling steps, our approach outperforms the previous state-of-the-art method (i.e., LiDARGen [73]) by a large margin for all considered metrics. Additionally, with only 50 evaluation steps, our method performs competitive with LiDARGen with a longer diffusion process of 1160 steps. Simultaneously, LiDMs reports improvements between $9.7\%$ and $31.0\%$ relative to the baseline model LDM [51]. For a qualitative comparison, in Fig. 3 we provide examples generated with each model, alongside reference point clouds. We further provide some 32-beam samples for example in Fig. 4.
@@ 5 Limitations
LiDM suffers from a visual gap relative to real-world LiDAR data. As shown in previous work [51], diffusion models should be powerful enough to capture the semantic information of input data, and therefore we argue that autoencoders are the key to failures when recovering most details of scenes. Our contributions are a step towards LiDAR-realistic autoencoders, however further work is still required. For example, autoencoders may reconstruct blurry boundaries between objects and the background, which though imperceptible on range images, may lead to visually unreasonable objects in 3D space.

### v2

@@ Abstract
Diffusion models (DMs) excel in photo-realistic image synthesis, but their adaptation to LiDAR scene generation poses a substantial hurdle. This is primarily because DMs operating in the point space struggle to preserve the curve-like patterns and 3D geometry of LiDAR scenes, which consumes much of their representation power. In this paper, we propose LiDAR Diffusion Models (LiDMs) to generate LiDAR-realistic scenes from a latent space tailored to capture the realism of LiDAR scenes by incorporating geometric priors into the learning pipeline. Our method targets three major desiderata: pattern realism, geometry realism, and object realism. Specifically, we introduce curve-wise compression to simulate real-world LiDAR patterns, point-wise coordinate supervision to learn scene geometry, and patch-wise encoding for a full 3D object context. With these three core designs, we establish a new state of the art on unconditional LiDAR generation in 64-beam scenario, while maintaining high efficiency compared to point-based DMs (up to 107$\times$ faster). Furthermore, by compressing LiDAR scenes into a latent space, we enable the controllability of DMs with various conditions such as semantic maps, camera views, and text prompts.
Figure 1: Our method (LiDM) establishes a new state-of-the-art in unconditional LiDAR-realistic scene generation, and marks a milestone towards conditional LiDAR scene generation from different input modalities.
@@ 1 Introduction
  • Our method achieves state-of-the-art on unconditional scene synthesis under 64-beam scenario, while realizing a $\times$107 speedup compared to point-based DMs.
@@ 2 Related Work
Given the larger-scale and complex scenes, treating LiDAR point clouds as a point cloud generation task akin to hand-crafted 3D models encounters difficulties. In [7], the authors explore the possibilities of generative models in LiDAR scenes by providing two solutions, LiDARVAE and LiDARGAN. [67] and [75] further introduce vector-quantized variational autoencoder and point-based diffusion model, respectively, to generate satisfactory samples of LiDAR scenes. However, these aforementioned methods still fail to generate LiDAR-realistic scenes as they may overlook the curve-like structures and geometric details inherent in LiDAR data.
@@ 4.2 Unconditional LiDAR Diffusion
Perceptual Statistical Method FRID $\downarrow$ FSVD $\downarrow$ FPVD $\downarrow$ JSD $\downarrow$ MMD $\downarrow$ ($10^{-4}$) Noise 3277 497.1 336.2 0.360 32.09 LiDARGAN [7] 1222 183.4 168.1 0.272 4.74 LiDARVAE [7] 199.1 129.9 105.8 0.237 7.07 ProjectedGAN [54] 149.7 44.7 33.4 0.188 2.88 UltraLiDAR [67] 370.0 72.1 66.6 0.747 17.12 LiDARGen [75] (1160s) 129.1 39.2 33.4 0.188 2.88 LiDARGen [75] (50s) 2051 480.6 400.7 0.506 9.91 LDM [51] (50s) 199.5 70.7 61.9 0.236 5.06 LiDM (ours, 50s) 158.8 53.7 42.7 0.213 4.46 $\Delta$ Improv. 20.4% 24.0% 31.0% 9.7% 11.9% LiDM (ours, 50s) 125.1 38.8 29.0 0.211 3.84 $\Delta$ Improv. 37.3% 45.1% 53.2% 10.6% 24.1%
Table 1: Comparison of unconditional LiDAR scene generation with recent state-of-the-art methods. We conduct experiments on 64-beam (i.e., KITTI-360 [37]) data.$\downarrow$” indicates that lower values are better. $N$-s refers to $N$ sampling steps during inference. $\Delta$ Improv. is the relative improvement of our method compared to the baseline of Latent Diffusion (LDM) [51], with the same number of diffusion steps. Note that, the diffusion process of LiDARGen [75] has 232 levels and 5 iterations in each level (i.e., 1160 steps in total).   denotes baseline results, while   /   denotes our results. $\dagger$: Exactly the same settings except the architecture of autoencoders. We evaluate each method with 2000 randomly generated samples.
To verify the effectiveness of our method, we train unconditional LiDMs on 64-beam data from KITTI-360 [37]. Following the practice of FID [21], we generate 2000 samples for the evaluation of unconditional generation. We evaluate the performance of each method using FRID (range image), FSVD (sparse volume), and FPVD (point-based volume), as described in Sec. 3.6. Together, these three metrics enable the perceptual evaluation to assess the quality of LiDAR scenes generated with different methods.
As shown in Table 1, with a very limited number of sampling steps (i.e., 50), we establish a new state of the art for almost all considered metrics. Specifically, within 50 sampling steps, our approach outperforms the previous state-of-the-art method (i.e., LiDARGen [75]) by a large margin for all considered metrics. Additionally, with only 50 evaluation steps, our method performs competitive with LiDARGen with a longer diffusion process of 1160 steps. LiDM reports $9.7\%\sim 31.0\%$ improvement over baseline model LDM [51] after replacing only the autoencoder. With all the techniques mentioned in Sec. 3.3, LiDM further improves by $10.6\%\sim 53.2\%$ over LDM. For a qualitative comparison, in Fig. 3 we provide examples generated with each model, alongside reference point clouds. We further provide some 32-beam samples for example in Fig. 4.
@@ 5 Limitations
Even though LiDM establishes a new state of the art, its generated samples still have a visual gap relative to real-world LiDAR data. As shown in previous work [51], diffusion models should be powerful enough to capture the semantic information of input data, and therefore we argue that autoencoders are the key to failures when recovering most details of scenes. Our contributions are a step towards LiDAR-realistic autoencoders, however further work is still required. For example, autoencoders may reconstruct blurry boundaries between objects and the background, which though imperceptible on range images, may lead to visually unreasonable objects in 3D space.